OpenStudy (anonymous):

Rose just bought a new jet ski for $5200, and it takes about $45.25 to fill the tank each time she goes out for a ride on the lake. Write a linear function that models these expenses. Also, interpret the slope and y-intercept, and use the line to calculate Rose’s total cost of the jet ski if she has gone out on 20 trips.

OpenStudy (anonymous):

a. y=5200x + 45.25 b. y=42.25-5200 c. y=-5200x-45.25 d. y=45.25x+5200

OpenStudy (anonymous):

this question is just a few questions up the list but without the y-intercept bit. the y-intercept is the initial cost = the cost borne by Rose without going out for ANY rides on the lake. to find the cost of 20 trips, use x = 20, which is R = 20 in the answer I posted in the previous incarnation of this question - above in the list.